OpenStudy (anonymous):

Help! How do I approach this question: Obtain an explicit formula for function(x) of f(x)=sum when n=1 to infinity of (nx^n)

OpenStudy (anonymous):

@Jhannybean

OpenStudy (anonymous):

Recall that for \(|x|<1\), you have \[\begin{align*}\sum_{n=0}^\infty x^n&=1+x+x^2+\cdots\\ &=\frac{1}{1-x} \end{align*}\] Suppose you take the derivative: \[\begin{align*}\sum_{n=0}^\infty nx^{n-1}&=\frac{d}{dx}\left[\frac{1}{1-x}\right]\\\\ &=\frac{1}{(1-x)^2} \end{align*}\] Multiply both sides by \(x\): \[\begin{align*}x\sum_{n=0}^\infty nx^{n-1}&=\sum_{n=0}^\infty nx^n\\\\ &=\sum_{n=1}^\infty nx^n \end{align*}\]

OpenStudy (mathmath333):

small correction there \[\frac{ d }{ dx }(\frac{ 1 }{ 1-x })=\frac{ 1 }{ (x-1)^2 }\]

OpenStudy (anonymous):

@mathmath333 we have the same result :) \[\frac{1}{(x-1)^2}=\frac{1}{[(-1)(1-x)]^2}=\frac{1}{(1-x)^2}\]

OpenStudy (mathmath333):

oh yes how can i miss that, lol

OpenStudy (anonymous):

Thank you guys!

OpenStudy (zarkon):

or let \[S=x+2x^2+3x^3+\cdots\] then \[xS=x^2+2x^3+3x^4+\cdots\] subtract \[S-xS=x+x^2+x^3+x^2+\cdots=\sum_{n=1}^{\infty}x^n=x\sum_{n=1}^{\infty}x^{n-1}=\frac{x}{1-x}\] \[S(1-x)=\frac{x}{1-x}\] \[S=\frac{x}{(1-x)^2}\]

OpenStudy (mathmath333):

cool

OpenStudy (zarkon):

no one pointed out my typo \[S-xS=x+x^2+x^3+x^{\color{red}{4}}+\cdots=\sum_{n=1}^{\infty}x^n=x\sum_{n=1}^{\infty}x^{n-1}=\frac{x}{1-x}\]